OpenStudy (anonymous):

Pendulum problem: An elephant has legs that stretch 2.3 m from its shoulders to the ground. How much time is required for one leg to swing forward after completing a stride?

OpenStudy (anonymous):

I tried the equations: \[2*\Pi*\sqrt{\frac{ L}{ g }}\] and \[2\pi \sqrt{\frac{ L }{ 3g }} \] but none of those seem to work.

OpenStudy (ksaimouli):

\[T = 2(\pi)\sqrt{(L/g) }\] this should work

OpenStudy (ksaimouli):

\[\frac{ 2\pi }{ \sqrt{\frac{ 2.3 }{ 9.8 }} }\]

OpenStudy (anonymous):

I figured it out. This problem represents a physical pendulum and has to factor in inertia of the leg. So the proper equation to use is \[\frac{ 1 }{ 2 \pi }\sqrt{\frac{ 3g }{ 2L }} \]